Home‎ > ‎Crew 133‎ > ‎

115-CommanderReport

Commander Report
01/15/2014
Paula Crock
 
Our project work is becoming focused on wrapping up and our thoughts are increasingly turning to our return journey.   Komorowski took us through an updated version of his telemedicine training today and has begun to pack up his gear.  Gartrelle and Diaz-Rubin did the final work with the lettuce seedlings, splitting up the root cubes to separate the plants in their hydroponics pods.  Jessup and Diaz-Rubin strung marker ribbons from the radio antenna support wires to make them more visible to new crew members when they are on EVA.  I’m working out a better way to attach the temperature monitoring devices to the EVA backpacks so that the incoming crews can more easily use them. Conversation around the Hab has become more about what we’ll do first when we get back to Earth and when we’ll visit one another, less about who is doing which report for CapComms or what is for dinner tomorrow or whether we have enough water for showers.
  
But not all of our projects and plans are winding down just yet.  We finally have some clear skies so Diaz-Rubin and Jessup have been getting some observing time in the past two nights.  Also, with the local area around the Hab mapped somewhat for the radio comms study, Howell and I took the ATVs out for the first time today to collect comms data some distance north of the Hab.  We were hampered by muddy roads so Gartrelle and Diaz-Rubin will attempt the EVA again tomorrow, starting out earlier in an attempt to complete the route before the roads become too muddy.

=====

CommanderCheckIn Report
01/15/2014
Paula Crock
Crew Physical Health:
All are well.
  
Eva Depart Time:
10:08
Eva Return Time:
13:16
Narrative Of Field Mission Results:
Cross road north of Hab was too muddy to traverse so planned route wasn't able to be completed but some radio comms and temperature data were collected
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Hab Maint:
N/A
Rep Sched:
Commander Check-In
Engineering Report
Science Report (GreenHab)
Science Report (Astro)
Future EVA Plan
EVA Report
Daily Photos
Journalist Report
Commander Report
Tomorrow Plan:
Retry of today's EVA since the cross road was too muddy by late morning when Howell/Crock attempted today
Inventory:
I am putting together a list of supplies that we have a sufficient supply of but the next crew will need to be provided with
Support Req:
Still waiting on info on lights/ballast for lab lights
Miscelleneous:
N/A